one year ago, ram's age was 8 times his son's age.Now his age is equal to square of his son's age .find their present age.
Answer by stanbon(75887)

one year ago, ram's age was 8 times his son's age.Now his age is equal to square of his son's age .find their present age.
-------
Equations:
r-1 = 8(s-1)
r = s^2
----
Substitute for "r" and solve for "s":
s^2-1 = 8s-8
s^2-8s+7 = 0
Factor:
(s-7)(s-1) = 0
Acceptable solution:
s = 7 yrs (son's age now)
----
r = s^2 = 49 (father's age now)
==============
Cheers,
Stan H.
===============